Find the derivative of y=x-4 and name the rules that you used.


4x+7x^2 all divided by x, find the derivative.


Find the values of x for y=-5x^2-2x
1
Expert's answer
2012-11-09T06:28:07-0500
y'=(x-4)'=x'-4'=1-0=1
(2x+7x^2)/x=2+7x
2'+7x'=7
y=-5x^2-2x
y=0 at x=0 and x=0.4